a pipe can fill a pool in 8 hours another pipe fill in 12 hours. if both pipes are applied to fill pool than how many hours it will take???
Answer by nerdybill(7384)

a pipe can fill a pool in 8 hours another pipe fill in 12 hours. if both pipes are applied to fill pool than how many hours it will take???
.
Let x = time (hrs) it takes for both
then
x(1/8 + 1/12)= 1
multiplying both sides by 96:
x(12 + 8)= 96
x(20)= 96
x = 4.8 hours
or
x = 4 hours and 48 minutes